OpenStudy (anonymous):

x^3+1 ? real and imagi. zeros

OpenStudy (zzr0ck3r):

try x=-1 what d you get?

OpenStudy (anonymous):

\[ (a+bi)^3 = -1 \]

OpenStudy (anonymous):

To better understand, think of it as: \[ (a+bi)^3 = -1+0i \]

OpenStudy (anonymous):

isn't the answer umm real=3 or 1 and for imagi= x=+-1?

OpenStudy (zzr0ck3r):

3^3+1 = 27+1=28 1^1+1 = 1+1 = 2

OpenStudy (anonymous):

Do you know complex analysis, or how to convert a number to polar form?

OpenStudy (zzr0ck3r):

-1 is a solution but it is not imaginary 1 is not a solution and is not imaginary (has no imaginary part...)

OpenStudy (anonymous):

y'all are thinking hard about this question

OpenStudy (anonymous):

I'm not even thinking about it. I already know the answer.

OpenStudy (anonymous):

The real challenge is getting you to understand.

OpenStudy (anonymous):

nvm i got it already y'alls answer is way off sorry but my teacher just helped me out

OpenStudy (anonymous):

I never gave you an answer.

OpenStudy (anonymous):

We didn't give you an answer. We we're proposing ways to solve this.

OpenStudy (anonymous):

lol ur formula confuses me

OpenStudy (zzr0ck3r):

I think that maybe you should go read some of the material leading up to this question. The fact that you said 3, was a solution of this, and that +-1 is an imaginary number, lets me know that you do not have the base material needed to complete this question.

OpenStudy (zzr0ck3r):

no insulting you, I cant do my homework either lol:)

OpenStudy (zzr0ck3r):

not*

OpenStudy (anonymous):

The only way I can think to solve this without knowing much about complex numbers is to discover that \(-1\) is a root, to then divided by \(x-(-1)\), to get a quadratic expression, and use the quadratic formula to get the remaining roots.

OpenStudy (anonymous):

i just did it already lol. thanks tho' and i messed up i looked at the wrong answer instead the real answer is real=3,1 and for imagi= 2, 1

OpenStudy (zzr0ck3r):

non of those numbers are solutions to the equation and the imaginary numbers you listed are not imaginary

OpenStudy (anonymous):

it's actually right tho' and plus my teacher was the one who helped me like few minutes ago and he said that that's the answer

OpenStudy (anonymous):

imagi 2,0 *

OpenStudy (zzr0ck3r):

ok

OpenStudy (zzr0ck3r):

good luck

OpenStudy (anonymous):

thanks ^^ i have a finals tom and i hope i can pass it